f(x)=|x²-x| differentiable at x=2. find its derivative

Grade:Upto college level

1 Answers

Sher Mohammad IIT Delhi
10 years ago
yes,
f(x)=x^2-x for x>1
f'(x)=2x-1
f'(2)=4-1=3
sher mohammad
b.tech, iit delhi